If the mean height of a group of students equals 170cm with a standard deviation of 10 cm, then the probability that a student is between 160cm and 180cm is .



1
Expert's answer
2021-11-15T04:16:31-0500

Mean = 170 cm

Standard deviation = 10 cm

"P(160<X<180)=P(\frac{160-170}{10}<Z<\frac{180-170}{10})"


"P(160<X<180)=P(-1<Z<1)"


From Z normal distribution table, it is found that


"P(160<x<180)=P(Z<1)-P(Z<-1)"


"P(160<x<180)=0.8413-0.1586=0.6827"